# Inferring Stellar Densities with Flexible Models I: The Distribution of RR Lyrae in the Milky Way with Gaia DR3

### v2

@@ Abstract
Understanding the formation and evolutionary history of the Milky Way requires detailed mapping of its stellar components, which preserve fossil records of the Galaxy’s assembly through cosmic time. RR Lyrae stars are particularly well-suited for this endeavor, as they are old, standard candle variables that probe the Galaxy’s earliest formation epochs. In this work, we employ a hierarchical Bayesian Gaussian Mixture Model (GMM) to characterize the three-dimensional density distribution of RR Lyrae stars in the Milky Way. This approach provides a flexible framework for modeling complex stellar distributions, particularly in the inner Galaxy where the bulge, disk, and halo components overlap. Our analysis reveals that the inner Galaxy is dominated by a distinct prolate stellar population with axis ratio qq=1.31. Consistent with previous work, we find the halo follows a $r^{-4}$ power-law profile that flattens within 12 kpc of the Galactic center. We also confirm the halo is oblate qq=0.70 with a tilt angle of $18^{\circ}$. We report for the first time that this tilt aligns the halo major axis in the direction of the Sagittarius dwarf galaxy. These results establish GMMs as an effective and flexible tool for modeling Galactic structure and provide new constraints on the distribution of old stars in the inner Galaxy.
@@
software: Astropy (Astropy Collaboration et al., 2013, 2018), Matplotlib (Hunter, 2007), IPython (Pérez & Granger, 2007), Numpy (Harris et al., 2020), Scipy (Virtanen et al., 2020), JAX (Bradbury et al., 2018), OTT-JAX (Cuturi et al., 2022), Optax (DeepMind et al., 2020), JAXopt (Blondel et al., 2021), BlackJAX (Cabezas et al., 2024), Corner (Foreman-Mackey, 2016) luas (Fortune et al., 2024)
1 Introduction
@@
With the billions of resolvable stars in the Milky Way, we can study the formation and evolutionary history of our Galaxy in exquisite detail. The Galaxy’s past structure is chronicled in the distribution of its resolved stellar populations (Grebel, 1999; Feast & Whitelock, 2000; Sandage & Tammann, 2006; Catchpole et al., 2016). By studying the structure of a stellar population with a constrained age distribution we obtain a glimpse into the Galaxy’s dynamical history since that population’s birth.
@@
Among the most valuable tracers for this purpose, RR Lyrae (RRL) are pulsating horizontal branch stars that are predominantly old ($\geq$10 Gyr) and metal-poor ($\rm{[Fe/H]\lesssim-0.5}$; Smith, 1995), making them an effective probe for the Galaxy’s earliest evolution. Furthermore, they are an ideal population for studying Galactic structure because their standard candle nature gives precise distance estimates as low as 1-2% when both infrared and optical photometry is available (Madore & Freedman, 2012; Neeley et al., 2017; Sesar et al., 2017a). Even in high extinction regions, RRL are still excellent distance indicators because they are color standards; they have the same effective temperature and therefore intrinsic color during their minimum light phase (Sturch, 1966; Guldenschuh et al., 2005; Kunder et al., 2010; Vivas et al., 2017).
@@
These exceptional properties have made RRL valuable for studying many aspects of the Galaxy’s structure, including the Galactic disk. The disk is typically modeled with two components: the older thick disk and the younger thin disk (Gilmore & Reid, 1983; Wyse & Gilmore, 1995; Bensby et al., 2011; Fuhrmann, 2008; Reddy et al., 2006). Given that they are generally old, RR Lyrae have been primarily used to study the Galactic thick disk (Layden, 1995; Amrose & Mckay, 2001; Kinemuchi et al., 2006; Mateu et al., 2012). Overall, the thick disk red giant branch (RGB) and red clump (RC) density distribution is consistent with results for RR Lyrae with most estimates converging around an exponential scale height ($h_{z}$) of $\sim$600-800 pc and scale length ($h_{R}$) of $\sim$ 2.1 kpc (Bovy et al., 2012a, b, 2016b; Robin et al., 2014; Mateu & Vivas, 2018; Tkachenko et al., 2025).
@@
The relationship between RRL and the thin disk, however, has proven more complex. While early studies suggest that the thin disk does not contain RRL (Martin & Morrison, 1998), recent work has revealed a small fraction of disk RRL with thin disk kinematics, sparking a debate about their origins (Chadid et al., 2017; Marsakov et al., 2018, 2019; Zinn et al., 2020; Prudil et al., 2020; Cabrera-Gadea et al., 2024). One hypothesis is that these stars have an accreted origin (Feuillet et al., 2022). Another possibility is that the thin disk RRL are actually an intermediate-age population that is a result of binary evolution (Pietrzyński et al., 2012; Karczmarek, 2012; Bobrick et al., 2024; Cuevas-Otahola et al., 2025).
@@
Beyond the Galactic disk, the metal-poor RRL serve as ideal tracers for the stellar density distribution of the Galactic halo, which is thought to consist of ancient accreted galaxies (De Lucia & Helmi, 2008; Deason et al., 2013). Generally, results for the halo density distribution of RRL are consistent with the total stellar distribution which are both most commonly modeled as a broken power law with a slope between $\alpha\approx$-1.5 to -3 for the inner halo and a slope of $\alpha\approx$-4 for the outer halo with the break radius around $\approx$20-30 kpc (Sesar et al., 2007; Watkins et al., 2009; Deason et al., 2011; Sesar et al., 2013; Faccioli et al., 2014; Zinn et al., 2014; Das & Binney, 2016; Hernitschek et al., 2018; Iorio et al., 2018; Han et al., 2022; Yang et al., 2022; Chen et al., 2023; Lane et al., 2023). None of these previous works include data within 5 kpc of the Galactic center with only a few of them including data with Galactoccetnric radii <10 kpc. There is also an interesting trend in the literature where works that include data at smaller radii tend to measure flatter slopes for the inner halo. Furthermore, previous works find that the halo is flattened with $q\approx$0.6-0.7 where $R=\sqrt{X^{2}+Y^{2}+(Z/q)^{2}}$ (Deason et al., 2011; Iorio et al., 2018; Han et al., 2022) with most recent work also measuring a tilt in the halo with respect to the Galactic disk plane (Iorio & Belokurov, 2019; Han et al., 2022).
@@
Perhaps the most challenging region to characterize, the inner Galaxy contains many overlapping components (halo, disk and bulge) and complex structures (the Galactic bar and X-shape). Whether the bulge is a distinct component from the inner halo is currently debated (Lucey et al., 2021; Ardern-Arentsen et al., 2024). Pietrukowicz et al. (2015) found the RRL density distribution in the inner Galaxy (0.2-3 kpc) is consistent with the measurements of the inner halo with a power law slope of $\alpha\approx$-3 (Pérez-Villegas et al., 2017). There is also evidence that a population of RR Lyrae in the inner Galaxy trace the bar structure (Kunder et al., 2020).
@@
In this work, we simultaneously model the RRL density distribution of the Milky Way bulge, halo and disk. In order to flexibly model all these structures, we utilize a Gaussian Mixture Model. Gaussian Mixtures have previously been used to model stellar density distributions of external galaxies, including to approximate power laws (Cappellari, 2002; Hogg & Lang, 2013; Miller & van Dokkum, 2021). However, this work is the first time this method has been applied to measure the stellar density distribution of the Milky Way.
@@
This publication is laid out as follows. In Section 2 we describe the sample of RR Lyrae, including parameterizing the selection function and calculating the distance modulus. We introduce our hierarchical Bayesian model of Gaussian mixtures in Section 3. We present our resulting density distribution of RRL and a comparison to previous work in Section 4. Last, in Section 5, we summarize this work and the conclusions.
2 RR Lyrae Sample
@@
Figure 1: The sky distribution in Galactic coordinates ($l,b$) of our RR Lyrae sample. stars in our sample. The top panel shows the log density for the entire sample of 128,353 RR Lyrae. After removing the XMCs with a cut in RA and DEC, there are 105,325 RRL that we use in our analysis as shown in the middle panel. The bottom panel shows the sky distribution of our final sample colored by the Gaia G magnitude, which, because of their standard-candle nature, is strongly correlated with the distance.
@@
Large samples of RR Lyrae stars are effective tools for studying Galactic structure given that their standard candle nature provides precise distance estimates. In this work, we use a catalog compiled from the three largest public surveys of RR Lyrae stars: Gaia DR3 Specific Objects Study (SOS, Clementini et al., 2023) ASAS-SN (Jayasinghe et al., 2019) and PanSTARRS1 (PS1, Sesar et al., 2017b). The quality cuts applied to each catalog are described in detail in Sec. 2 of Mateu et al. (2020), as well as the cross-matching strategy used to match the known RR Lyrae from each catalog to a source in the Gaia DR3 gaia_source table. Duplicate sources were removed based on Gaia’s source_id, resulting in a base catalog with a total of 288,928 RR Lyrae.
@@
We present the sky distribution of our sample in Galactic coordinates ($l,b$) in Figure 1. The top panel shows the log density of RRL for the entire sample. The LMC and SMC appear as two blobs in the region $30^{\circ}<l<90^{\circ}$ and $b\approx-30^{\circ}$. As these are major structures in the sample which we do not include as part of the Milky Way’s halo, we remove them from our sample. The simplest way to remove them from the sample without complicated selection function effects from distance uncertainties is a cut in right ascension (RA) and declination (DEC). We remove stars with RA<110 and DEC <-55. The sky distribution of the resulting 105,325 RRL is shown in the middle panel of Figure 1.
@@
The bottom panels of Figure 1 shows the sky distribution of our final sample colored by the observed G magnitude. The Sagittarius dwarf galaxy stream is visible as a streak of stars with high G magnitudes passing vertically through the Galactic center at $(l,b)=(0,0)^{\circ}.$ However, as this is much lower density than the XMCs we do not find it necessary to remove from the sample. Along the disk plane at $b=0^{\circ}$, high levels of dust extinction cause higher G magnitudes as well as gaps in the RR Lyrae distribution. We account for this effect in our selection function which is described in the following Section.
2.1 Selection Function
@@
Figure 2: Empirical Completeness Maps of the RR Lyrae sample in Galactic Coordinates at distances of 5 kpc (top left), 10 kpc (middle left), and 20 kpc (bottom left). We show corresponding slices of the data in distance bins of 1 kpc in the right panels. At small distances, the completeness is degraded because of the bright limits of surveys. The dominant structure in the completeness map is consistent with the Gaia scanning law.
@@
In order to account for the observational biases of our RR Lyrae sample, we need to quantify the selection function by estimating the completeness of our sample as a function of sky position and distance. We estimate the completeness following the procedure described in detail in Mateu et al. (2020) and introduced by Rybizki & Drimmel (2018). The method uses two independent catalogs, in our case Gaia and ASAS+PS1, to empirically determine the completeness for each catalog based on the number of stars in both catalogs as well as in each. In Mateu et al. (2020) completeness maps were presented for the Gaia DR2 SOS, ASAS-SN and PS1 RR Lyrae catalogs, and an update presented in Mateu (2024) for Gaia DR3 and the combined Gaia+ASAS+PS1. In the latter, completeness maps using only the RUWE and phot_bp_rp_excess_factor (BEP) filters were reported. Here we compute the completeness (or selection function) maps including also the more stringent availability filter on the Li et al. (2023) photometric metallicity (or equivalently Gaia DR3’s). The resulting selection function maps for RRab and RRc stars with RUWE, BEP and photometric metallicities filters are publicly available at the rrl_completeness GitHub repository [^1] https://github.com/cmateu/rrl_completeness.
@@
In the left panels of Figure 2 we present images of the resulting selection function in Galactic coordinates evaluated at three distances (5, 10, and 20 kpc from top to bottom). The right panels show corresponding 1 kpc slices of the data. The dominant structure is a result of the Gaia scanning law (see Figure 10 in Cantat-Gaudin et al., 2023) where the completeness is lowest in regions with the fewest scans. At small distances, the completeness is lower due to the bright limits of the surveys, while the faint limits impact the completeness at large distances.
3 Hierarchical Bayesian Model of Multi-Gaussian Galaxy

### v3

@@ Abstract
Understanding the formation and evolutionary history of the Milky Way requires detailed mapping of its stellar components, which preserve fossil records of the Galaxy’s assembly through cosmic time. RR Lyrae stars are particularly well-suited for this endeavor, as they are old, standard candle variables that probe the Galaxy’s earliest formation epochs. In this work, we employ a hierarchical Bayesian Gaussian Mixture Model (GMM) to characterize the three-dimensional density distribution of RR Lyrae stars in the Milky Way over the galactocentric radius ($R$) of $\approx 0.2-120~\rm{kpc}$. This approach provides a flexible framework for modeling complex stellar distributions, particularly in the inner Galaxy where the bulge, disk, and halo components overlap. Our analysis reveals that the inner Galaxy ($R\lesssim 10~\rm{kpc}$) is dominated by a distinct prolate stellar population with axis ratio qq=1.31. Consistent with previous work, we find the halo follows a $r^{-4}$ power-law profile that flattens within 12 kpc of the Galactic center. We also confirm the outer halo ($R\gtrsim 10~\rm{kpc}$) is oblate qq=0.70 with a tilt angle of $18^{\circ}$. We report for the first time that this tilt aligns the halo major axis in the direction of the Sagittarius dwarf galaxy. These results establish GMMs as an effective and flexible tool for modeling Galactic structure and provide new constraints on the distribution of old stars in the inner Galaxy.
@@
software: Astropy (Astropy Collaboration et al., 2013, 2018), Matplotlib (Hunter, 2007), IPython (Pérez and Granger, 2007), Numpy (Harris et al., 2020), Scipy (Virtanen et al., 2020), JAX (Bradbury et al., 2018), OTT-JAX (Cuturi et al., 2022), Optax (DeepMind et al., 2020), JAXopt (Blondel et al., 2021), BlackJAX (Cabezas et al., 2024), Corner (Foreman-Mackey, 2016) luas (Fortune et al., 2024)
I Introduction
@@
With the billions of resolvable stars in the Milky Way, we can study the formation and evolutionary history of our Galaxy in exquisite detail. The Galaxy’s past structure is chronicled in the distribution of its resolved stellar populations (Grebel, 1999; Feast and Whitelock, 2000; Freeman and Bland-Hawthorn, 2002; Sandage and Tammann, 2006; Catchpole et al., 2016). By studying the structure of a stellar population with a constrained age distribution we obtain a glimpse into the Galaxy’s dynamical history since that population’s birth. As different galaxy formation processes lead to different correlations between stellar age and dynamics, we can use these observed patterns to distinguish between formation scenarios (Eggen et al., 1962; Searle and Zinn, 1978; Freeman, 1987; Molla et al., 1996; Walker et al., 1996; Bland-Hawthorn and Gerhard, 2016; Helmi, 2020; Deason and Belokurov, 2024).
@@
Among the most valuable tracers for this purpose, RR Lyrae (RRL) are pulsating horizontal branch stars that are predominantly old ($\geq$10 Gyr) and metal-poor ($\rm{[Fe/H]\lesssim-0.5}$; Smith, 1995), making them an effective probe for the Galaxy’s earliest evolution. Furthermore, they are an ideal population for studying Galactic structure because their standard candle nature gives precise distance estimates as low as 1-2% when both infrared and optical photometry is available (Madore and Freedman, 2012; Neeley et al., 2017; Sesar et al., 2017a; Mullen et al., 2023). Even in high extinction regions, RRL are still excellent distance indicators because they are color standards; they have the same effective temperature and therefore intrinsic color during their minimum light phase (Sturch, 1966; Guldenschuh et al., 2005; Kunder et al., 2010; Vivas et al., 2017).
@@
These exceptional properties have made RRL valuable for studying many aspects of the Galaxy’s structure, including the Galactic disk. The disk is typically modeled with two components: the older thick disk and the younger thin disk (Gilmore and Reid, 1983; Wyse and Gilmore, 1995; Bensby et al., 2011; Fuhrmann, 2008; Reddy et al., 2006). Given that they are generally old, RRL have been primarily used to study the Galactic thick disk (Layden, 1995; Amrose and Mckay, 2001; Kinemuchi et al., 2006; Mateu et al., 2012). Overall, the thick disk red giant branch (RGB) and red clump (RC) density distribution is consistent with results for RRL with most estimates converging around an exponential scale height ($h_{z}$) of $\approx 600-800$ pc and scale length ($h_{R}$) of $\approx 2.1$ kpc (Bovy et al., 2012a, b, 2016b; Robin et al., 2014; Mateu and Vivas, 2018; Tkachenko et al., 2025).
@@
The relationship between RRL and the thin disk, however, has proven more complex. Although, metal-rich RRL stars with thin disk kinematics were found by Layden (1995) and as early as Strugnell et al. (1986), other studies were unable to confirm the association of these stars to the thin disk and suggested they might belong to the metal-rich tail of the thick disk (Martin and Morrison, 1998). Recent work has revealed a significant fraction of disk RRL with thin disk kinematics (Iorio and Belokurov, 2021), sparking a debate about their origins (Chadid et al., 2017; Marsakov et al., 2018, 2019; Zinn et al., 2020; Prudil et al., 2020). One hypothesis is that these stars have an accreted origin (Feuillet et al., 2022). Another possibility is that the thin disk RRL are actually an intermediate-age population and that it could be the result of binary evolution (Iorio and Belokurov, 2021; Bobrick et al., 2024; Cuevas-Otahola et al., 2025; Cabrera-Gadea et al., 2024; Zhāng et al., 2025).
@@
Beyond the Galactic disk, the metal-poor RRL serve as ideal tracers for the stellar density distribution of the Galactic halo, which is thought to consist of ancient accreted galaxies (De Lucia and Helmi, 2008; Deason et al., 2013). Generally, results for the halo density distribution of RRL are consistent with the total stellar distribution which are both most commonly modeled as a broken power law with a slope between $\alpha\approx-1.5$ to 3-3 for the inner halo and a slope of $\alpha\approx-4$ for the outer halo with the break radius around $\approx 20-30$ kpc (Sesar et al., 2007; Watkins et al., 2009; Deason et al., 2011; Sesar et al., 2013; Faccioli et al., 2014; Zinn et al., 2014; Das and Binney, 2016; Hernitschek et al., 2018; Iorio et al., 2018; Chen et al., 2023; Lane et al., 2023). However, other works have found doubly broken power-law slopes with the second break radius of $\approx 10~\rm{kpc}$ (Han et al., 2022; Yang et al., 2022). Recent works have also found the slope may be distinct along different lines-of-sight in the Galaxy (Hernitschek et al., 2018; Amarante et al., 2024; Medina et al., 2024). Works focusing on the very distant outer halo ($R\gtrsim 100~\rm{kpc}$) have found similar slopes of $\alpha\approx-4$ (Thomas et al., 2018; Medina et al., 2018; Stringer et al., 2021; Feng et al., 2024), although some have found a break radius at $R\approx 160-210~\rm{kpc}$ (Fukushima et al., 2018, 2019). None of these previous works include data within 5 kpc of the Galactic center with only a few of them including data with Galactocentric radii <10 kpc. There is also an interesting trend in the literature where works that include data at smaller radii tend to measure flatter slopes for the inner halo. Furthermore, previous works find that the halo is flattened with $q\approx$0.6-0.7 where $R=\sqrt{X^{2}+Y^{2}+(Z/q)^{2}}$ (Deason et al., 2011; Iorio et al., 2018; Han et al., 2022) with most recent work also measuring a tilt in the halo with respect to the Galactic disk plane (Iorio and Belokurov, 2019; Han et al., 2022). There is also work which finds the flattening varies with Galactocentric radius (Xue et al., 2015; Xu et al., 2018; Iorio et al., 2018; Hernitschek et al., 2018).
@@
Perhaps the most challenging region to characterize, the inner Galaxy contains many overlapping components (halo, disk and bulge) and complex structures (the Galactic bar and X-shape). Whether the bulge is a distinct component from the inner halo is currently debated (Lucey et al., 2021; Ardern-Arentsen et al., 2024). Pietrukowicz et al. (2015) found the RRL density distribution in the inner Galaxy (0.2-3 kpc) is consistent with the measurements of the inner halo with a power law slope of $\alpha\approx$-3 (Pérez-Villegas et al., 2017). There is also evidence that a population of RRL in the inner Galaxy trace the bar structure (Kunder et al., 2020).
@@
In this work, we simultaneously model the RRL density distribution of the Milky Way bulge, halo and disk. In order to flexibly model all these structures, we utilize a Gaussian Mixture Model (GMM). GMMs have previously been used to model stellar density distributions of external galaxies, including to approximate power laws (Cappellari, 2002; Hogg and Lang, 2013; Miller and van Dokkum, 2021). However, this work is the first time this method has been applied to measure the stellar density distribution of the Milky Way.
@@
This publication is laid out as follows. In Section II we describe the sample of RRL, including parameterizing the selection function and calculating the distance modulus. We introduce our hierarchical Bayesian model of Gaussian mixtures in Section III. We present our resulting density distribution of RRL and a comparison to previous work in Section IV. Last, in Section V, we summarize this work and the conclusions.
II RR Lyrae Sample
@@
Figure 1: The sky distribution in Galactic coordinates ($l,b$) of our RRL sample. The top panel shows the log density for the 105,325 RRL we use in this work, after removing the LMC and SMC with a cut in RA and DEC. The bottom panel shows the sky distribution of our final sample colored by the Gaia G magnitude, which, because of their standard-candle nature, is strongly correlated with the distance.
@@
Large samples of RRL stars are effective tools for studying Galactic structure given that their standard candle nature provides precise distance estimates. In this work, we use a catalog compiled from the three largest public surveys of RRL stars: Gaia DR3 Specific Objects Study (SOS, Clementini et al., 2023) ASAS-SN (Jayasinghe et al., 2019) and PanSTARRS1 (PS1, Sesar et al., 2017b). The quality cuts applied to each catalog are described in detail in Sec. 2 of Mateu et al. (2020), as well as the cross-matching strategy used to match the known RRL from each catalog to a source in the Gaia DR3 gaia_source table. Duplicate sources were removed based on Gaia’s source_id, resulting in a base catalog with a total of 288,928 RRL.
@@
We present the sky distribution of our sample in Galactic coordinates ($l,b$) in Figure 1. To remove contamination from the two most massive dwarf galaxies orbiting the Milky Way, the LMC and SMC, we remove stars with $\mathrm{RA}<110^{\circ}$ and $\mathrm{Dec.}<-55^{\circ}$. This is the simplest way to remove these major structures from the sample without complicated selection function effects from distance uncertainties. The sky distribution of the resulting 102,761 RRL is shown in the top panel of Figure 1.
@@
The bottom panel of Figure 1 shows the sky distribution of our final sample colored by the observed G magnitude. The Sagittarius dwarf galaxy stream is visible as a streak of stars with high G magnitudes passing vertically through the Galactic center at $(l,b)=(0,0)^{\circ}.$ However, as this is much lower density than the LMC and SMC we do not find it necessary to remove from the sample. Along the disk plane at $b=0^{\circ}$, high levels of dust extinction cause higher G magnitudes as well as gaps in the RRL distribution. We account for this effect in our selection function which is described in the following Section.
II.1 Selection Function
@@
Figure 2: Empirical Completeness Maps of the RRL sample in Galactic Coordinates at heliocentric distances of 5 kpc (top left), 10 kpc (middle left), and 20 kpc (bottom left). We show corresponding slices of the data in heliocentric distance bins of 1 kpc in the right panels. At small distances, the completeness is degraded because of the bright limits of surveys. The dominant structure in the completeness map is consistent with the Gaia scanning law.
@@
In order to account for the observational biases of our RRL sample, we need to quantify the selection function by estimating the completeness of our sample as a function of sky position and heliocentric distance. We estimate the completeness following the procedure described in detail in Mateu et al. (2020) and introduced by Rybizki and Drimmel (2018). The method uses two independent catalogs, in our case Gaia and ASAS+PS1, to empirically determine the completeness for each catalog based on the number of stars in both catalogs as well as in each. Completeness is estimated separately for RRL of types ab and c [^1] For these purposes we consider the very few RRd as RRab since they are too few for a completeness map to be estimated., since RRc samples are well known to be more affected by incompleteness as these stars have lower photometric amplitudes and are more prone to misclassification (see e.g. Mateu et al., 2020; Sesar et al., 2017b). In Mateu et al. (2020) completeness maps were presented for the Gaia DR2 SOS, ASAS-SN and PS1 RRL catalogs, and an update presented in Mateu (2024) for Gaia DR3 and the combined Gaia+ASAS+PS1. In the latter, completeness maps using only the RUWE and phot_bp_rp_excess_factor (BEP) filters were reported. Here we compute the completeness (or selection function) maps including also the more stringent availability filter on the Li et al. (2023) photometric metallicity (or equivalently Gaia DR3’s). The resulting selection function maps with RUWE, BEP and photometric metallicities filters are publicly available in Mateu (2025). The tools to generate this and other similar selection functions are also available at the rrl_completeness GitHub repository [^2] https://github.com/cmateu/rrl_completeness .
@@
In the left panels of Figure 2 we present images of the resulting selection function in Galactic coordinates evaluated at three heliocentric distances (5, 10, and 20 kpc from top to bottom). The right panels show corresponding 1 kpc slices of the data. The dominant structure is a result of the Gaia scanning law (see Figure 10 in Cantat-Gaudin et al., 2023) where the completeness is lowest in regions with the fewest scans. At small heliocentric distances, the completeness is lower due to the bright limits of the surveys, while the faint limits impact the completeness at large heliocentric distances.
III Hierarchical Bayesian Model of Multi-Gaussian Galaxy